# CLIP-HandID: Vision-Language Model for Hand-Based Person Identification

### v2

@@ I Introduction
Recently, vision-language cross-modal learning such as Contrastive Language-Image Pre-training (CLIP) [10], which establishes meaningful connections between visual representations and their textual descriptions, has received great attention in deep learning. This model employs two key innovations: training on extensive datasets (image-text pairs) and reformulating the pre-training objective to align visual features with their linguistic counterparts using self-supervised contrastive learning, which minimizes cosine distance (maximizes cosine similarity) for matched image-text pairs. This paradigm enables the image encoder to capture richer semantic information form images with the help of textual prompts through text encoder and develop highly adaptable features that demonstrate strong transferability across diverse uni-modal and multi-modal downstream tasks [11, 12, 13], when compared to uni-modal CNN-based [14] and Vision Transformer (ViT)-based [15] methods pre-trained on manually labelled images (ImageNet) from a set of predefined classes where visual information conveying semantic meaning beyond the predefined categories is totally disregarded. Consider a standard image classification scenario: the predefined class labels serve as concrete textual inputs that can be incorporated into descriptive prompts (e.g., ”A photo of a [class]”, where [class] is generally replaced by concrete text labels). The classification process occurs through cross-modal comparison - matching the visual features extracted from images through the image encoder against the textual features generated by processing these label-enhanced prompts through the text encoder - as a zero-shot solution during inference. This vision-language model has been adopted for person and/or vehicle re-identification in [16, 17] where it has shown promising results. However, there is no work which has leveraged the vision-language foundation models such as CLIP for identification of a person based on hand images for criminal investigations. In fine-grained hand-based person identification, the labels are indexes (numeric values) instead of text descriptions. Therefore, we adopted textual inversion [18, 19] to describe the images with specific words to fully exploit the cross-modal description ability in CLIP for the development of hand-based person identification. This textual inversion network learns pseudo-tokens to generate customized textual descriptions for individual subjects. These pseudo-tokens enable the model to focus on semantically or meaningfully relevant visual regions that align precisely with the prompt context. This design allows the model to learn fine-grained attribute information.
@@ II Proposed Method
where $K$ is the total number of classes (identities) and $p_{k}$ is the predicted probability of class $k$. We also use label smoothing [22] which prevents over-fitting and over-confidence with smoothing value ($\epsilon$) of 0.1. The ground-truth distribution over labels $q_{k}$ by including label smoothing can be given as
@@
The image encoder $\theta_{I}(.)$ takes an image $\mathbf{x}_{i}\in\mathbb{R}^{H\times W\times C}$ as input, where H, W and C are the height, the width and the number of channels, respectively. Similarly, the text encoder $\theta_{T}(.)$ takes a tokenized textual prompt $\mathbf{t}_{i}\in\mathbb{R}^{N\times D}$ as input, where N is the number of text sequences and D is the context length - the maximum number of tokens that the model can process in a single input sequence. Thus, the image feature $I_{i}=\theta_{I}(\mathbf{x}_{i})$ and text feature $T_{i}=\theta_{T}(\mathbf{t}_{i})$. Hence, to calculate supervised contrastive loss, we first need to compute the similarity between the embeddings of input image $\mathbf{x}_{i}$ and input tokenized textual prompt $\mathbf{t}_{i}$ as follows
@@
where $I_{i}=\theta_{I}(\mathbf{x}_{i})$ and $T_{i}=\theta_{T}(\mathbf{t}_{i})$ are the image feature and the text feature, respectively, projected into a cross-modal embedding space, and $\text{sim}(u,v)=\frac{u\cdot v}{\|u\|_{2}\|v\|_{2}}$ is the cosine similarity. Then, the image-to-text contrastive loss $\mathcal{L}_{i2t}$ is calculated as:
@@
where the numerators in (4) and (5) represent the similarity scores between matched embedding pairs, while the denominators aggregate all similarity measures relative to the anchor $I_{i}$ or $T_{i}$, and $\tau$ represents the temperature, which we set $\tau=1$.
@@ III-A Experimental Setup
Implementation details: Our implementation of CLIP-HandID was developed using the PyTorch framework, with training performed on an NVIDIA GeForce RTX 2080 Ti GPU. As a backbone, we use both ResNet-50 and ViT-B/16 pre-trained from CLIP as our image encoder and a pre-trained text encoder (CLIP text Transformer). In addition, the CLIP-HandID incorporates the inversion network which is random-initialized. During training, input images were first resized to $256\times 256$ pixels, then randomly cropped to $224\times 224$. We applied several augmentation techniques including random horizontal flip, normalization, and color jittering. To mitigate over-fitting and excessive confidence in predictions, we employed label smoothing [22] with a smoothing factor ($\epsilon$) of 0.1. For testing, images were simply resized to $224\times 224$ and normalized, without random cropping or other augmentation. In both training and testing phases, we randomized images order through dataset reshuffling. The model was trained for 70 epochs with a batch size of 20 for ViT-B/16 and 4 for ResNet-50 backbones, using cross-entropy and supervised contrastive losses and the Adam optimizer. We applied L2 regularization with a weight decay factor of $5\times 10^{-4}$. A learning rate warmup strategy [23] was implemented for the first 10 epochs, gradually increasing the rate from $5\times 10^{-8}$ to $5\times 10^{-6}$. Subsequent learning rate adjustments reduced it to $2.5\times 10^{-6}$ at epoch 40 and $1.25\times 10^{-6}$ at epoch 60. Notably, we assigned different learning rates to various network components: the backbone network’s existing layers used one-tenth the learning rate of newly added layers (inversion network, FC or linear layers and batch normalizations), with proper initialization of weights and biases.

### v3

@@ I Introduction
Recently, vision-language cross-modal learning such as Contrastive Language-Image Pre-training (CLIP) [10], which establishes meaningful connections between visual representations and their textual descriptions, has received great attention in deep learning. This model employs two key innovations: training on extensive datasets (image-text pairs) and reformulating the pre-training objective to align visual features with their linguistic counterparts using self-supervised contrastive learning, which minimizes cosine distance (maximizes cosine similarity) for matched image-text pairs. This paradigm enables the image encoder to capture richer semantic information from images with the help of textual prompts through text encoder and develop highly adaptable features that demonstrate strong transferability across diverse uni-modal and multi-modal downstream tasks [11, 12, 13], when compared to uni-modal CNN-based [14] and Vision Transformer (ViT)-based [15] methods pre-trained on manually labelled images (ImageNet) from a set of predefined classes where visual information conveying semantic meaning beyond the predefined categories is totally disregarded. Consider a standard image classification scenario: the predefined class labels serve as concrete textual inputs that can be incorporated into descriptive prompts (e.g., ”A photo of a [class]”, where [class] is generally replaced by concrete text labels). The classification process occurs through cross-modal comparison - matching the visual features extracted from images through the image encoder against the textual features generated by processing these label-enhanced prompts through the text encoder - as a zero-shot solution during inference. This vision-language model has been adopted for person and/or vehicle re-identification in [16, 17] where it has shown promising results. However, there is no work which has leveraged the vision-language foundation models such as CLIP for identification of a person based on hand images for criminal investigations. In fine-grained hand-based person identification, the labels are indexes (numeric values) instead of text descriptions. Therefore, we adopted textual inversion [18, 19] to describe the images with specific words to fully exploit the cross-modal description ability in CLIP for the development of hand-based person identification. This textual inversion network learns pseudo-tokens to generate customized textual descriptions for individual subjects. These pseudo-tokens enable the model to focus on semantically or meaningfully relevant visual regions that align precisely with the prompt context. This design allows the model to learn fine-grained attribute information.
@@ II Proposed Method
where $K$ is the total number of classes (identities) and $p_{k}$ is the predicted probability of class (identity) $k$. We also use label smoothing [22] which prevents over-fitting and over-confidence with smoothing value ($\epsilon$) of 0.1. The ground-truth distribution over labels $q_{k}$ by including label smoothing can be given as
@@
The image encoder $\theta_{I}(.)$ takes an image $\mathbf{x}_{i}\in\mathbb{R}^{C\times H\times W}$ as input, where C, H and W are the number of channels, the height and the width, respectively. Similarly, the text encoder $\theta_{T}(.)$ takes a tokenized textual prompt $\mathbf{t}_{i}\in\mathbb{R}^{N\times D}$ as input, where N is the number of text sequences and D is the context length - the maximum number of tokens that the model can process in a single input sequence. Thus, the image feature $I_{i}=\theta_{I}(\mathbf{x}_{i})$ and text feature $T_{i}=\theta_{T}(\mathbf{t}_{i})$. Hence, to calculate supervised contrastive loss, we first need to compute the similarity between the embeddings of input image $\mathbf{x}_{i}$ and input tokenized textual prompt $\mathbf{t}_{i}$ as follows
@@
where $I_{i}=\theta_{I}(\mathbf{x}_{i})$ and $T_{i}=\theta_{T}(\mathbf{t}_{i})$ are the image feature and the text feature, respectively, projected into a cross-modal embedding space, and $\text{sim}(\mathbf{u},\mathbf{v})=\frac{\mathbf{u}\cdot\mathbf{v}}{\|\mathbf{u}\|_{2}\|\mathbf{v}\|_{2}}$ is the cosine similarity. Note that $\mathbf{u}\cdot\mathbf{v}=\mathbf{u}^{T}\mathbf{v}$. Then, the image-to-text contrastive loss $\mathcal{L}_{i2t}$ is calculated as:
@@
where the numerators in (4) and (5) represent the similarity scores between matched embedding pairs, while the denominators aggregate all similarity measures relative to the anchor $I_{i}$ or $T_{i}$, and $\tau$ represents the temperature, which we set $\tau=1$. Note that $I_{i}$ and $T_{i}$ have the same dimension in the cross-modal embedding space.
@@ III-A Experimental Setup
Implementation details: Our implementation of CLIP-HandID was developed using the PyTorch framework, with training performed on an NVIDIA GeForce RTX 2080 Ti GPU. As a backbone, we use both ResNet-50 and ViT-B/16 pre-trained from CLIP as our image encoder and a pre-trained text encoder (CLIP text Transformer). In addition, the CLIP-HandID incorporates the inversion network which is random-initialized. During training, input images were first resized to $\times 256$ 256 pixels, then randomly cropped to $\times 224$ 224. We applied several augmentation techniques including random horizontal flip, normalization, and color jittering. To mitigate over-fitting and excessive confidence in predictions, we employed label smoothing [22] with a smoothing factor ($\epsilon$) of 0.1. For testing, images were simply resized to $\times 224$ 224 and normalized, without random cropping or other augmentation. In both training and testing phases, we randomized images order through dataset reshuffling. The model was trained for 70 epochs with a batch size of 20 for ViT-B/16 and 4 for ResNet-50 backbones, using cross-entropy and supervised contrastive losses and the Adam optimizer. We applied L2 regularization with a weight decay factor of $5\times 10^{-4}$. A learning rate warmup strategy [23] was implemented for the first 10 epochs, gradually increasing the rate from $5\times 10^{-8}$ to $5\times 10^{-6}$. Subsequent learning rate adjustments reduced it to $2.5\times 10^{-6}$ at epoch 40 and $1.25\times 10^{-6}$ at epoch 60. Notably, we assigned different learning rates to various network components: the backbone network’s existing layers used one-tenth the learning rate of newly added layers (inversion network, FC or linear layers and batch normalizations), with proper initialization of weights and biases.